OpenStudy (anonymous):

Hi can someone please help me with a limit of a sequence? Lim as N approaches infinity {Tan^-1 n / n }

OpenStudy (anonymous):

\[ - \frac \pi 2< Tan^{-1} n< \frac \pi 2 \] What can you conclude?

OpenStudy (anonymous):

Hence \[ \left |\frac {\tan^{-1}(n)}{n} \right |< \frac {\pi } {2n} \] Can you finish it now?

OpenStudy (anonymous):

Im sorry I really dont see how you got from the first step to the second.....

OpenStudy (anonymous):

If a number x is between -5 and 5 then the absolute value of x is less than 5

OpenStudy (anonymous):

and if \[ | x| < 5 \] then \[ \left |\frac x n \right |< \frac 5 n \]